OpenStudy (anonymous):

state the domain and range f(x)=2(1/3)^x

OpenStudy (anonymous):

domain all of real range is positive R

OpenStudy (anonymous):

howd you get that?

OpenStudy (anonymous):

all the possible values of x which satisfies the relation will give you the domain.. n for every x the outcome gives you the range

OpenStudy (anonymous):

Basically you're looking for disallowed values for x. If for example you had \(\sqrt{x}\) for your function you would know that negative numbers for x are not allowed. If you had \(\frac{4}{x}\) then 0 would not be allowed. If you had \(\sqrt{5-x}\) then your domain would be numbers less than or equal to 5. etc

OpenStudy (anonymous):

shouldnt they be numbers?

OpenStudy (anonymous):

The domain and range are sets of numbers. There are many values you can plug in for x and get many different values for the function.

OpenStudy (anonymous):

so why is the domain and range not numbers?

OpenStudy (anonymous):

it is a set of numbers

OpenStudy (anonymous):

????

OpenStudy (anonymous):

you dont understand what is meant by a "set of numbers"??

OpenStudy (anonymous):

i just dont understand what your saying

OpenStudy (anonymous):

{1, 2, 3.4, 15} is a set of 4 numbers. (-5,1) is a set of infinitely many numbers between -5 and 1, but not including -5 or 1

OpenStudy (anonymous):

i get that!!! I just dont understand the domain and range

OpenStudy (anonymous):

Well the domain is just a set of numbers that are 'legal' to plug in for x.

OpenStudy (anonymous):

If your function is well behaved (doesn't divide or take the square root) then anything is legal. Otherwise you might have to restrict what x can be. If you divide by x, then x cannot be 0. If you divide by x+3 then x cannot be -3. If you take the square root of x, then x cannot be negative, etc.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Can you take the square root of a negative number?

OpenStudy (anonymous):

(and get a real result)

OpenStudy (anonymous):

I have to go, but we can talk about it another time if you like.

OpenStudy (anonymous):

margaret are you aware of complex numbers?

OpenStudy (anonymous):

no im not of complex numbers